# Spectral Representation-based Reinforcement Learning

### v1

@@ Abstract
In real-world applications with large state and action spaces, reinforcement learning (RL) typically employs function approximations to represent core components like the policies, value functions, and dynamics models. Although powerful approximations such as neural networks offer great expressiveness, they often present theoretical ambiguities, suffer from optimization instability and exploration difficulty, and incur substantial computational costs in practice. In this paper, we introduce the perspective of spectral representations as a solution to address these difficulties in RL. Stemming from the spectral decomposition of the transition operator, this framework yields an effective abstraction of the system dynamics for subsequent policy optimization while also providing a clear theoretical characterization. We reveal how to construct spectral representations for transition operators that possess latent variable structures or energy-based structures, which implies different learning methods to extract spectral representations from data. Notably, each of these learning methods realizes an effective RL algorithm under this framework. We also provably extend this spectral view to partially observable MDPs. Finally, we validate these algorithms on over 20 challenging tasks from the DeepMind Control Suite, where they achieve performances comparable or superior to current state-of-the-art model-free and model-based baselines.
Keywords: reinforcement learning, optimal control, representation learning, self-supervised learning, world models
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful tool in real-world applications that involve sequential decision-making (degrave2022magnetic; ouyang2022training; tang2025deep). It formulates tasks based on Markov Decision Processes (MDPs) and learns an optimal policy that can maximize the expected cumulative return through interacting with an unknown environment (sutton1998reinforcement). For MDPs with finite state and action spaces, known as the tabular case, efficient and convergent algorithms with well-defined theoretical sample complexities have been developed (auer2008near; dann2015sample; agrawal2017optimistic; azar2017minimax; jin2018q). However, the optimization cost of these algorithms quickly becomes unacceptable as the volume of the state space and action space grows beyond tabular cases, even towards infinity in continuous state-action cases. To tackle such scalability challenges, modern RL algorithms employ function approximations to parameterize the core components in the pipeline (jin2020provably; yang2020function; long2021perturbational). Based on which component is approximated, these algorithms can be roughly classified into two categories: model-free RL and model-based RL. Model-free methods directly parameterize the policy and/or value functions, which are then optimized in an end-to-end fashion with trial-and-error interactions with the environment (dai2018sbeed; fujimoto2018addressing; haarnoja2018soft). In contrast, model-based RL trains a surrogate dynamics model to capture the environment’s state transition, and the optimal policy can then be derived by planning or model predictive control using this learned model (hafner2019dream; janner2019trust).
The architectural simplicity of modern model-free RL has enabled its successful integration with powerful function approximations (e.g., deep neural networks) for policy/value function parameterization, leading to notable applications in diverse domains, such as generative model fine-tuning (ouyang2022training; yang2024using), video games (mnih2013playing; berner2019dota), and robotic control (tang2025deep). However, such expressive power also comes with theoretical challenges: with general function approximations, standard RL methods such as temporal-difference learning with off-policy data can diverge under some circumstances, a phenomenon known as the “deadly triad” (sutton1998reinforcement). Furthermore, model-free RL is completely blind to the underlying dynamics, ultimately leading to a prohibitive sample complexity in practice (fujimoto2018addressing; haarnoja2018soft).
@@
Model-based RL, on the other hand, explicitly learns a model of the environment’s dynamics and leverages this model to facilitate policy optimization (wang2019benchmarking; luo2024survey). When the model possesses a certain structure (e.g., linear quadratic regulators (LQRs)), it can be shown that optimal policies can be derived efficiently. However, in complex domains where the state transition can be nonlinear or non-deterministic, advanced modeling techniques (e.g., recurrent (hafner2019learning; hafner2019dream), attentional (chen2022transdreamer; micheli2022transformers), and diffusion-based models (ding2024diffusion; alonso2024diffusion)) with larger capacity become necessary in order to accurately reflect the true dynamics and reduce the approximation error. This introduces a dilemma analogous to that faced in model-free RL: with non-linear models, solving for an optimal policy via planning is computationally intractable, and becoming a long-standing challenging problem in control community (khalil2015nonlinear). Consequently, practical model-based RL must resort to approximate inference methods, such as Dyna-style planning (kurutach2018model; janner2019trust; luo2024survey) or backpropagating the gradient through the models (deisenroth2011pilco; heess2015learning). These compromises not only risk converging to suboptimal solutions, but also incur substantial costs both during training and inference due to their intricate designs, and thus waste the flexibility in modeling and effort in learning.
@@
Motivated by the recent success of representation learning, we present spectral representations as an affirmative answer to the above question. Specifically, spectral representations stem from the functional decomposition of the transition probability operator in MDP, thereby encoding sufficient information about the underlying dynamics. Leveraging an analysis analogous to the linear MDP literature (jin2020provably), it can be demonstrated that the $Q$-value functions in RL can be sufficiently expressed with spectral representations, which paves the way for efficient learning and planning. Although several algorithms have been developed to leverage the power of spectral representations for reinforcement learning (ren2022free; zhang2022making; ren2023latent; ren2023spectral; zhang2023provable; shribak2024diffusion), this field nevertheless lacks a coherent framework to unify these methods, elucidate their connections, and conduct fair comparison. This survey aims to fill this blank by organizing these disparate methods into a cohesive overview and systematically benchmarking them through controlled experiments.
We illustrate the structure of this survey in Figure 1. Specifically, after introducing the background knowledge in Section 2, Section 3 begins with a comprehensive overview of the spectral representation, covering its definitions and properties (Section 3.1) and its construction with different dynamics structures (Section 3.2). To address the intractability of spectral decomposition and maximum likelihood estimation (MLE) in large-scale spaces, Section 4 reviews practical algorithms for learning these representations directly from data. Next, Section 5 details how to integrate these learned representations into the reinforcement learning pipeline. Section 6 extends the framework to partially observable Markov decision processes (POMDPs), thereby accommodating more realistic decision-making scenarios. Finally, to validate the effectiveness as well as to facilitate further research in this direction, Section 7 presents an extensive empirical study. We release a unified codebase [^1] Code is available at https://github.com/spectral-rl/spectral-rl. and benchmark performance on over 20 continuous control tasks from the DeepMind Control Suite (tassa2018deepmind), encompassing diverse settings that include both fully observable state-based inputs and partially observable visual data. Generally, the algorithms based on spectral representations outperform their model-free counterparts, especially as task dynamics become more complex. In visual settings, these methods achieve performance comparable or superior to state-of-the-art model-based algorithms while obviating the need for computationally expensive trajectory synthesis. Finally, through a detailed ablation analysis, we investigate the influence of certain key design choices to provide a deeper understanding of these methods.
@@ 3.1 Spectral View of MDPs
While seminal work in reinforcement learning has proposed efficient algorithms for tabular MDPs with finite state and action spaces (auer2008near; dann2015sample; agrawal2017optimistic), scaling to large or continuous spaces has remained a challenge. Recent works have turned to function approximations to overcome such challenges. A prominent approach in this direction is the linear (low-rank) MDP hypothesis (jin2020provably), which posits that the transition operator admits a low-rank linear decomposition. Inspired by this, we consider a more generalized version without the finite rank assumption. Specifically, for any well-behaved transition operator $\mathbb{P}$ and reward function $r$, there always exist two feature maps $\bm{\phi}:{\mathcal{S}}\times\mathcal{A}\to\mathcal{H}$, $\bm{\mu}:{\mathcal{S}}\to\mathcal{H}$ and a vector $\bm{\theta}_{r}\in\mathcal{H}$ for some proper Hilbert space $\mathcal{H}$, such that for $\forall(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have
@@
In fact, since the feature maps and the vector can be infinite-dimensional, such a decomposition always exists for the transition kernel $\mathbb{P}$. To see this, the singular value decomposition (SVD) of the transition operator (mollenhauer2020singular) gives a concrete example:
@@
A key consequence of this linear structure is that the $Q$-value functions of any given policy $\pi$ also inherit this linearity (yao2014pseudo; jin2020provably; yang2020reinforcement).
@@
Proof 3.2.
The linearity of $Q$-value functions can be examined by the Bellman equation:
@@
To better situate the proposed spectral representations in the vast representation learning literature, we summarize existing representations that stem from spectral decomposition of transition-related matrices in Table 1. As discussed, spectral representations are derived from the left singular vectors of the single-step transition operator. In contrast, successor features (dayan1993improving; kulkarni2016deep; barreto2017successor) are actually the left singular vector of $(I-\gamma\mathbb{P}^{\pi})^{-1}$, as proved in (ren2023spectral), which summarizes future state visitation from given states under a prefixed policy. The Laplacian-based methods (wu2018laplacian) compute representations from the eigenvectors of the symmetrized transition matrix with the largest eigenvalues. Finally, if the reward function is known a priori, the Krylov basis (petrik2007analysis) suggests using the series $\{(\mathbb{P}^{\pi})^{i}r\}_{i=1}^{k}$ as the representation for value functions.
@@
Successor Feature (dayan1993improving) $\texttt{svd}\left((I-\gamma\mathbb{P}^{\pi})^{-1}\right)$
Laplacian (wu2018laplacian) $\texttt{eig}\left(\Lambda\mathbb{P}^{\pi}+(\mathbb{P}^{\pi})^{\top}\Lambda\right)$
Krylov Basis (petrik2007analysis) $\{(\mathbb{P}^{\pi})^{i}r\}_{i=1}^{k}$
@@
Exploiting this structure, hu2024primal reduces the saddle-point optimization of the off-policy evaluation problem to a convex-concave objective, using $\bm{\phi}(s,a)$ as the primal representation for $Q$-value functions and $\bm{\mu}^{\pi}(s,a)$ as the dual representation for the occupancy measure.
@@ 3.2.1 Linear Formulation
By comparing this linear formulation (11) with (6), we instantly identify the spectral representation as $\bm{\phi}(s,a)=\bm{\varphi}(s,a)\in\mathbb{R}^{d}$. This is theoretically equivalent to the linear MDP hypothesis in jin2020provably; however, we note that this structure is limited, as the finite-rank linearity assumption is strong and restricts the model’s expressive capacity.
@@ 3.2.2 Latent Variable Formulation
where $\varphi(\cdot|s,a):{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta\left(\mathcal{Z}\right)$ and $\nu\left(\cdot|z\right):\mathcal{Z}\rightarrow\Delta\left({\mathcal{S}}\right)$. Notably, if the measures $\varphi(\cdot|s,a)\in L_{2}$ and $\nu(s^{\prime}|\cdot)\in L_{2}$, the integral in the latent variable formulation can also be formulated as an inner product (ren2023latent):
@@
In fact, this structure is widely adopted in model-based reinforcement learning and world models with varied specific choices for the latent space and distributions. For instance, setting $\mathcal{Z}=\mathbb{R}^{d}$ and using Gaussian distributions for $p$ recovers the dynamics modeling used in PlaNet (hafner2019learning) and Dreamer (hafner2019dream); while setting $Z=\{1,2,\ldots,D\}$ and $p(\cdot|s,a)$ as Categorical distributions recovers the models in DreamerV2 (hafner2020mastering) and DreamerV3 (hafner2023mastering). Moreover, such latent dynamics can also be built upon learned state representations, as demonstrated by zhou2411dino and xiang2025pan.
@@
Since $\varphi(\cdot|s,a)$ must be a valid probability measure, latent variable models are thus limited in expressiveness as compared to the unconstrained version. In fact, agarwal2020flambe has demonstrated that the simplex feature $\bm{\phi}_{\text{LV}}$ can be exponentially weaker than the linear counterpart $\bm{\phi}_{\text{Linear}}$, meaning that $\bm{\phi}_{\text{LV}}$ may require exponentially more dimensions than $\bm{\phi}_{\text{Linear}}$ to characterize the same transition operator. However, as the theoretical characterization in Section 5 suggests, the apparent dimension may not be a good measure of the complexity of the representations. As long as the transition operator satisfies some regularity conditions, sampling-efficient learning is still possible.
@@ 3.2.3 Energy-based Formulation
This final expression represents the transition as an inner product in $\mathcal{H}_{k}$, the reproducing kernel Hilbert space of the Gaussian kernel $k$, and satisfies the decomposition in (6). However, this RKHS representation $k(\bm{\varphi}(s,a),\cdot)$ and $k(\cdot,\bm{\nu}(s^{\prime}))$ is conceptual and implicit. To derive a concrete representation, ren2022free and shribak2024diffusion adopt random Fourier features (rahimi2007random), as defined in the following lemma.
Lemma 3.3.
(Random Fourier features (rahimi2007random)) Consider a shift invariant kernel $k(\bm{x},\bm{y})=k(\bm{x}-\bm{y})$ on $\mathbb{R}^{d}$. Define $\zeta_{\bm{\omega}}(\bm{x})=e^{j\bm{\omega}^{\top}\bm{x}}$, we have
@@
The case of stochastic nonlinear control (zheng2022optimistic; ren2022free; ren2025stochastic; ma2025offline), whose state-space dynamics follows the form $s^{\prime}=\bm{\varphi}(s,a)+v$ where $v$ is a Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$, can be understood as a special instance of either the energy-based formulation or latent-variable formulation. In this case, the state transition dynamics can be written as $\mathbb{P}(s^{\prime}|s,a)\propto\exp\left(-\frac{\|s^{\prime}-\bm{\varphi}(s,a)\|^{2}}{2\sigma^{2}}\right)$. By applying the same random Fourier features, we can recover the spectral structure analogous to (22). Alternatively, the connection to the latent-variable formulation can be seen through the decomposition:

### v2

@@ Abstract
In real-world applications with large state and action spaces, reinforcement learning (RL) typically employs function approximations to represent core components like the policies, value functions, and dynamics models. Although powerful approximations such as neural networks offer great expressiveness, they often present theoretical ambiguities, suffer from optimization instability and exploration difficulty, and incur substantial computational costs in practice. In this paper, we introduce the perspective of spectral representations as a solution to address these difficulties in RL. Stemming from the spectral decomposition of the transition operator, this framework yields an effective abstraction of the system dynamics for subsequent policy optimization while also providing a clear theoretical characterization. We reveal how to construct spectral representations for transition operators that possess latent variable structures or energy-based structures, which implies different learning methods to extract spectral representations from data. Notably, each of these learning methods realizes an effective RL algorithm under this framework. We also provably extend this spectral view to partially observable MDPs. Finally, we validate these algorithms on over 20 challenging tasks from the DeepMind Control Suite, where they achieve performances comparable or superior to current state-of-the-art model-free and model-based baselines. Our code is publicly released at https://spectral-rl.github.io.
Keywords— reinforcement learning, optimal control, representation learning, self-supervised learning, world models
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful tool in real-world applications that involve sequential decision-making (Degrave et al., 2022; Ouyang et al., 2022; Tang et al., 2025). It formulates tasks based on Markov Decision Processes (MDPs) and learns an optimal policy that can maximize the expected cumulative return through interacting with an unknown environment (Sutton et al., 1998). For MDPs with finite state and action spaces, known as the tabular case, efficient and convergent algorithms with well-defined theoretical sample complexities have been developed (Auer et al., 2008; Dann and Brunskill, 2015; Agrawal and Jia, 2017; Azar et al., 2017; Jin et al., 2018). However, the optimization cost of these algorithms quickly becomes unacceptable as the volume of the state space and action space grows beyond tabular cases, even towards infinity in continuous state-action cases. To tackle such scalability challenges, modern RL algorithms employ function approximations to parameterize the core components in the pipeline (Jin et al., 2020; Yang et al., 2020; Long and Han, 2021). Based on which component is approximated, these algorithms can be roughly classified into two categories: model-free RL and model-based RL. Model-free methods directly parameterize the policy and/or value functions, which are then optimized in an end-to-end fashion with trial-and-error interactions with the environment (Dai et al., 2018; Fujimoto et al., 2018; Haarnoja et al., 2018). In contrast, model-based RL trains a surrogate dynamics model to capture the environment’s state transition, and the optimal policy can then be derived by planning or model predictive control using this learned model (Hafner et al., 2019a; Janner et al., 2019).
The architectural simplicity of modern model-free RL has enabled its successful integration with powerful function approximations (e.g., deep neural networks) for policy/value function parameterization, leading to notable applications in diverse domains, such as generative model fine-tuning (Ouyang et al., 2022; Yang et al., 2024), video games (Mnih et al., 2013; Berner et al., 2019), and robotic control (Tang et al., 2025). However, such expressive power also comes with theoretical challenges: with general function approximations, standard RL methods such as temporal-difference learning with off-policy data can diverge under some circumstances, a phenomenon known as the “deadly triad” (Sutton et al., 1998). Furthermore, model-free RL is completely blind to the underlying dynamics, ultimately leading to a prohibitive sample complexity in practice (Fujimoto et al., 2018; Haarnoja et al., 2018).
@@
Model-based RL, on the other hand, explicitly learns a model of the environment’s dynamics and leverages this model to facilitate policy optimization (Wang et al., 2019; Luo et al., 2024). When the model possesses a certain structure (e.g., linear quadratic regulators (LQRs)), it can be shown that optimal policies can be derived efficiently. However, in complex domains where the state transition can be nonlinear or non-deterministic, advanced modeling techniques (e.g., recurrent (Hafner et al., 2019b, a), attentional (Chen et al., 2022a; Micheli et al., 2022), and diffusion-based models (Ding et al., 2024; Alonso et al., 2024)) with larger capacity become necessary in order to accurately reflect the true dynamics and reduce the approximation error. This introduces a dilemma analogous to that faced in model-free RL: with non-linear models, solving for an optimal policy via planning is computationally intractable, and becoming a long-standing challenging problem in control community (Khalil and Minor, 2015). Consequently, practical model-based RL must resort to approximate inference methods, such as Dyna-style planning (Kurutach et al., 2018; Janner et al., 2019; Luo et al., 2024) or backpropagating the gradient through the models (Deisenroth and Rasmussen, 2011; Heess et al., 2015). These compromises not only risk converging to suboptimal solutions, but also incur substantial costs both during training and inference due to their intricate designs, and thus waste the flexibility in modeling and effort in learning.
@@
Motivated by the recent success of representation learning, we present spectral representations as an affirmative answer to the above question. Specifically, spectral representations stem from the functional decomposition of the transition probability operator in MDP, thereby encoding sufficient information about the underlying dynamics. Leveraging an analysis analogous to the linear MDP literature (Jin et al., 2020), it can be demonstrated that the $Q$-value functions in RL can be sufficiently expressed with spectral representations, which paves the way for efficient learning and planning. Although several algorithms have been developed to leverage the power of spectral representations for reinforcement learning (Ren et al., 2022; Zhang et al., 2022; Ren et al., 2023a, b; Zhang et al., ; Shribak et al., 2024), this field nevertheless lacks a coherent framework to unify these methods, elucidate their connections, and conduct fair comparison. This survey aims to fill this blank by organizing these disparate methods into a cohesive overview and systematically benchmarking them through controlled experiments.
We illustrate the structure of this survey in Figure 1. Specifically, after introducing the background knowledge in Section 2, Section 3 begins with a comprehensive overview of the spectral representation, covering its definitions and properties (Section 3.1) and its construction with different dynamics structures (Section 3.2). To address the intractability of spectral decomposition and maximum likelihood estimation (MLE) in large-scale spaces, Section 4 reviews practical algorithms for learning these representations directly from data. Next, Section 5 details how to integrate these learned representations into the reinforcement learning pipeline. Section 6 extends the framework to partially observable Markov decision processes (POMDPs), thereby accommodating more realistic decision-making scenarios. Finally, to validate the effectiveness as well as to facilitate further research in this direction, Section 7 presents an extensive empirical study. We release a unified codebase and benchmark performance on over 20 continuous control tasks from the DeepMind Control Suite (Tassa et al., 2018), encompassing diverse settings that include both fully observable state-based inputs and partially observable visual data. Generally, the algorithms based on spectral representations outperform their model-free counterparts, especially as task dynamics become more complex. In visual settings, these methods achieve performance comparable or superior to state-of-the-art model-based algorithms while obviating the need for computationally expensive trajectory synthesis. Finally, through a detailed ablation analysis, we investigate the influence of certain key design choices to provide a deeper understanding of these methods.
@@ 3.1 Spectral View of MDPs
While seminal work in reinforcement learning has proposed efficient algorithms for tabular MDPs with finite state and action spaces (Auer et al., 2008; Dann and Brunskill, 2015; Agrawal and Jia, 2017), scaling to large or continuous spaces has remained a challenge. Recent works have turned to function approximations to overcome such challenges. A prominent approach in this direction is the linear (low-rank) MDP hypothesis (Jin et al., 2020), which posits that the transition operator admits a low-rank linear decomposition. Inspired by this, we consider a more generalized version without the finite rank assumption. Specifically, for any well-behaved transition operator $\mathbb{P}$ and reward function $r$, there always exist two feature maps $\bm{\phi}:{\mathcal{S}}\times\mathcal{A}\to\mathcal{H}$, $\bm{\mu}:{\mathcal{S}}\to\mathcal{H}$ and a vector $\bm{\theta}_{r}\in\mathcal{H}$ for some proper Hilbert space $\mathcal{H}$, such that for $\forall(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have
@@
In fact, since the feature maps and the vector can be infinite-dimensional, such a decomposition always exists for the transition kernel $\mathbb{P}$. To see this, the singular value decomposition (SVD) of the transition operator (Mollenhauer et al., 2020) gives a concrete example:
@@
A key consequence of this linear structure is that the $Q$-value functions of any given policy $\pi$ also inherit this linearity (Yao et al., 2014; Jin et al., 2020; Yang and Wang, 2020).
@@
Proof The linearity of $Q$-value functions can be examined by the Bellman equation:
@@
To better situate the proposed spectral representations in the vast representation learning literature, we summarize existing representations that stem from spectral decomposition of transition-related matrices in Table 1. As discussed, spectral representations are derived from the left singular vectors of the single-step transition operator. In contrast, successor features (Dayan, 1993; Kulkarni et al., 2016; Barreto et al., 2017) are actually the left singular vector of $(I-\gamma\mathbb{P}^{\pi})^{-1}$, as proved in (Ren et al., 2023b), which summarizes future state visitation from given states under a prefixed policy. The Laplacian-based methods (Wu et al., 2018) compute representations from the eigenvectors of the symmetrized transition matrix with the largest eigenvalues. Finally, if the reward function is known a priori, the Krylov basis (Petrik, 2007) suggests using the series $\{(\mathbb{P}^{\pi})^{i}r\}_{i=1}^{k}$ as the representation for value functions.
@@
Successor Feature (Dayan, 1993) $\texttt{svd}\left((I-\gamma\mathbb{P}^{\pi})^{-1}\right)$
Laplacian (Wu et al., 2018) $\texttt{eig}\left(\Lambda\mathbb{P}^{\pi}+(\mathbb{P}^{\pi})^{\top}\Lambda\right)$
Krylov Basis (Petrik, 2007) $\{(\mathbb{P}^{\pi})^{i}r\}_{i=1}^{k}$
@@
Exploiting this structure, Hu et al. (2024) reduces the saddle-point optimization of the off-policy evaluation problem to a convex-concave objective, using $\bm{\phi}(s,a)$ as the primal representation for $Q$-value functions and $\bm{\mu}^{\pi}(s,a)$ as the dual representation for the occupancy measure.
@@ 3.2.1 Linear Formulation
By comparing this linear formulation (11) with (6), we instantly identify the spectral representation as $\bm{\phi}(s,a)=\bm{\varphi}(s,a)\in\mathbb{R}^{d}$. This is theoretically equivalent to the linear MDP hypothesis in Jin et al. (2020); however, we note that this structure is limited, as the finite-rank linearity assumption is strong and restricts the model’s expressive capacity.
@@ 3.2.2 Latent Variable Formulation
where $\varphi(\cdot|s,a):{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta\left(\mathcal{Z}\right)$ and $\nu\left(\cdot|z\right):\mathcal{Z}\rightarrow\Delta\left({\mathcal{S}}\right)$. Notably, if the measures $\varphi(\cdot|s,a)\in L_{2}$ and $\nu(s^{\prime}|\cdot)\in L_{2}$, the integral in the latent variable formulation can also be formulated as an inner product (Ren et al., 2023a):
@@
In fact, this structure is widely adopted in model-based reinforcement learning and world models with varied specific choices for the latent space and distributions. For instance, setting $\mathcal{Z}=\mathbb{R}^{d}$ and using Gaussian distributions for $p$ recovers the dynamics modeling used in PlaNet (Hafner et al., 2019b) and Dreamer (Hafner et al., 2019a); while setting $Z=\{1,2,\ldots,D\}$ and $p(\cdot|s,a)$ as Categorical distributions recovers the models in DreamerV2 (Hafner et al., 2020) and DreamerV3 (Hafner et al., 2023). Moreover, such latent dynamics can also be built upon learned state representations, as demonstrated by Zhou et al. and Xiang et al. (2025).
@@
Since $\varphi(\cdot|s,a)$ must be a valid probability measure, latent variable models are thus limited in expressiveness as compared to the unconstrained version. In fact, Agarwal et al. (2020) has demonstrated that the simplex feature $\bm{\phi}_{\text{LV}}$ can be exponentially weaker than the linear counterpart $\bm{\phi}_{\text{Linear}}$, meaning that $\bm{\phi}_{\text{LV}}$ may require exponentially more dimensions than $\bm{\phi}_{\text{Linear}}$ to characterize the same transition operator. However, as the theoretical characterization in Section 5 suggests, the apparent dimension may not be a good measure of the complexity of the representations. As long as the transition operator satisfies some regularity conditions, sampling-efficient learning is still possible.
@@ 3.2.3 Energy-based Formulation
This final expression represents the transition as an inner product in $\mathcal{H}_{k}$, the reproducing kernel Hilbert space of the Gaussian kernel $k$, and satisfies the decomposition in (6). However, this RKHS representation $k(\bm{\varphi}(s,a),\cdot)$ and $k(\cdot,\bm{\nu}(s^{\prime}))$ is conceptual and implicit. To derive a concrete representation, Ren et al. (2022) and Shribak et al. (2024) adopt random Fourier features (Rahimi and Recht, 2007), as defined in the following lemma.
Lemma 2
(Random Fourier features (Rahimi and Recht, 2007)) Consider a shift invariant kernel $k(\bm{x},\bm{y})=k(\bm{x}-\bm{y})$ on $\mathbb{R}^{d}$. Define $\zeta_{\bm{\omega}}(\bm{x})=e^{j\bm{\omega}^{\top}\bm{x}}$, we have
@@
The case of stochastic nonlinear control (Zheng et al., 2022; Ren et al., 2022, 2025; Ma et al., 2025), whose state-space dynamics follows the form $s^{\prime}=\bm{\varphi}(s,a)+v$ where $v$ is a Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$, can be understood as a special instance of either the energy-based formulation or latent-variable formulation. In this case, the state transition dynamics can be written as $\mathbb{P}(s^{\prime}|s,a)\propto\exp\left(-\frac{\|s^{\prime}-\bm{\varphi}(s,a)\|^{2}}{2\sigma^{2}}\right)$. By applying the same random Fourier features, we can recover the spectral structure analogous to (22). Alternatively, the connection to the latent-variable formulation can be seen through the decomposition: